# Armadillo: An Efficient Framework for Numerical Linear Algebra

### v1

@@ I Introduction
To address the above issues in a coherent framework, we have implemented the Armadillo linear algebra library for C++ [17], which automatically optimises mathematical expressions (both at compile-time and run-time) and efficiently maps them to BLAS/LAPACK routines, all while providing a user-friendly Matlab-like programming interface directly in C++. Armadillo essentially acts as a high-level domain specific language  [22, 15] built on top of the host C++ language, allowing for resource efficient numerical linear algebra without the many pain points of low-level code. This enables rapid and low risk conversion of research code into production environments, and even permits direct prototyping of algorithms within C++.
@@
Armadillo employs two strategies for automatically optimising mathematical expressions, both aiming to reduce computational effort: (i) compile-time fusion of operations to reduce the need for temporary objects, (ii) mixture of compile-time detection of expressions and run-time analysis of matrix properties, with the aim of re-ordering and translating operations. Both strategies extensively use C++ template metaprogramming concepts [23], where the compiler is induced to reason at compile-time to generate code tailored for each expression.
@@ II Expression Optimisation via Metaprogramming
The above pattern is detected at compile-time, and is automatically translated as a call to the xGESV function in LAPACK, which solves a system of linear equations without the matrix inverse operation.
@@
Analysis of matrix properties is also exploited in the evaluation of matrix multiplication chains. For example, in the expression ${\boldsymbol{A}}{\cdot}{\boldsymbol{B}}{\cdot}{\boldsymbol{C}}{\cdot}{\boldsymbol{D}}$, each of the possible matrix pairs is examined. The pair which results in the smallest matrix is executed first, thereby reducing computational effort in subsequent matrix multiplications. As such, it is possible for the entire expression to be evaluated right-to-left, rather than the traditional left-to-right order.
@@ III Empirical Evaluation
    ${\boldsymbol{E}}={\boldsymbol{A}}_{m{\times}m}\cdot{\boldsymbol{B}}_{m{\times}\frac{m}{2}}\cdot{\boldsymbol{C}}_{\frac{m}{2}{\times}\frac{m}{3}}\cdot{\boldsymbol{D}}_{\frac{m}{3}{\times}\frac{m}{4}}$. this is an instance of chained matrix multiplication resulting in a matrix. Here the matrices are progressively decreasing in size. A naive implementation would evaluate each of the matrix products in the standard left-to-right manner, disregarding the wider context of the expression. An optimised implementation can examine the sizes of all possible matrix products within the expression, and determine that evaluating the products in a reversed right-to-left order will save computational effort.
@@ IV Conclusion
As we have seen, Armadillo facilitates easy and maintainable representation of arbitrary linear algebra expressions in C++ that are efficiently mapped to underlying BLAS and LAPACK operations. Users do not need to worry about cumbersome manual memory management or complicated calls to BLAS and LAPACK subroutines. There is no performance penalty for the abstractions provided by Armadillo; in fact, as we have shown, in some cases Armadillo is able to avoid BLAS and LAPACK calls entirely via compile-time optimisations through template metaprogramming, and in many cases Armadillo provides the best solution to the linear algebra mapping problem [3, 16]. Armadillo can achieve considerable reductions in processing time over direct and/or naive implementations.
Work on Armadillo started in 2008. Over the years the library and has been iteratively and collaboratively developed with feedback from the wider scientific and engineering communities. The library is currently comprised of about 135,000 lines of templated code, excluding BLAS and LAPACK routines. Support is provided for matrices with single- and double-precision floating point elements (in both real and complex forms), as well as integer elements. Dense and sparse storage formats are supported.

### v2

@@ I Introduction
To address the above issues in a coherent framework, we have implemented the Armadillo linear algebra library for C++ [18], which automatically optimises mathematical expressions (both at compile-time and run-time) and efficiently maps them to BLAS/LAPACK routines, all while providing a user-friendly Matlab-like programming interface directly in C++. Armadillo essentially acts as a high-level domain specific language [16] built on top of the host C++ language, allowing for resource efficient numerical linear algebra without the many pain points of low-level code. This enables rapid and low risk conversion of research code into production environments, and even permits direct prototyping of algorithms within C++.
@@
Armadillo employs two strategies for automatically optimising mathematical expressions, both aiming to reduce computational effort: (i) compile-time fusion of operations to reduce the need for temporary objects, (ii) mixture of compile-time detection of expressions and run-time analysis of matrix properties, with the aim of re-ordering and translating operations. Both strategies extensively use C++ template metaprogramming concepts [9, 23], where the compiler is induced to reason at compile-time to generate code tailored for each expression.
@@ II Expression Optimisation via Metaprogramming
The above pattern is detected at compile-time, and is automatically translated as a call to the xGESV function in LAPACK, which solves a system of linear equations without the matrix inverse.
@@
Analysis of matrix properties is also exploited in the evaluation of matrix multiplication chains. For example, in the expression ${\boldsymbol{A}}{\cdot}{\boldsymbol{B}}{\cdot}{\boldsymbol{C}}{\cdot}{\boldsymbol{D}}$, each of the possible matrix pairs is examined. The pair which results in the smallest matrix is multiplied first, thereby reducing computational effort in subsequent matrix multiplications. As such, it is possible for the entire expression to be evaluated right-to-left (while respecting general non-commutativity of matrix multiplication), rather than the traditional left-to-right order.
@@ III Empirical Evaluation
    ${\boldsymbol{E}}={\boldsymbol{A}}_{m{\times}m}\cdot{\boldsymbol{B}}_{m{\times}\frac{m}{2}}\cdot{\boldsymbol{C}}_{\frac{m}{2}{\times}\frac{m}{3}}\cdot{\boldsymbol{D}}_{\frac{m}{3}{\times}\frac{m}{4}}$. this is an instance of chained matrix multiplication resulting in a matrix. Here the matrices are progressively decreasing in size. A naive implementation would evaluate each of the matrix products in the standard left-to-right manner, disregarding the wider context of the expression. An optimised implementation can examine the sizes of all possible matrix products within the expression, and determine that evaluating the products in a reversed order will save computational effort.
@@ IV Conclusion
As we have seen, Armadillo facilitates easy and maintainable representation of arbitrary linear algebra expressions in C++ that are efficiently mapped to underlying BLAS and LAPACK operations. Users do not need to worry about cumbersome manual memory management or complicated calls to BLAS and LAPACK subroutines. There is virtually no performance penalty for the abstractions provided by Armadillo. Moreover, through under-the-hood template metaprogramming and automatic optimisations of expressions, Armadillo can achieve considerable reductions in processing time over direct and/or naive implementations.
Work on Armadillo started in 2008. Over the years the library has been iteratively and collaboratively developed with feedback from the wider scientific and engineering communities. The library provides over 200 functions; in addition to elementary operations, there are functions for statistics, signal processing, non-contiguous submatrix views, and various matrix factorisations. The library is currently comprised of about 135,000 lines of templated code, excluding BLAS and LAPACK routines. Support is provided for matrices with single- and double-precision floating point elements (in both real and complex forms), as well as integer elements. Dense and sparse storage formats are supported.